### v1

@@ Appendix 0.B Additional Experiments
0.B.1 More Ablation Study
@@
Figure 8: More ablation study on Concept Enhancement Constraints (in Eq. 7) and Concept Isolation Constraints (in Eq. 8). The upper portion displays outcomes utilizing our full methodology, while the lower portion illustrates results with specific modules omitted, highlighting the significance of each component within our approach. The red boxes and the yellow boxes are used to accentuate the distinctions between the anime style and the real-world style, respectively.
@@
In the main ablation study Sec. 4.5, we explored the synergistic effects of combining modules with similar functionalities. Here, we delve deeper with an extensive ablation study on our Concept Enhancement Constraints, as outlined in Eq. 7 (which includes Gaussian sample strategy and $\mathcal{L}_{fill}$) and Concept Isolation Constraints (incorporating the concept region mask and $\mathcal{L}_{region}$ in Eq. 8). These examinations aim to illuminate their contributions to model performance, as illustrated in Fig. 8. Specifically, in the first column, the absence of Gaussian sampling leads to the concepts not being accurately centered within their designated boxes. This lack of precision can even cause anime concepts to appear outside their intended boundaries, resulting in a loss of their unique identity traits. In the second column, without $\mathcal{L}_{fill}$, both anime and realistic figures fail to occupy their designated boxes completely, pointing to a deficiency in fully utilizing the allocated space. In the third column, we observe the issue of concept confusion, such as the merging of anime haircuts with traits of realistic figures, which indicates a loss of distinctiveness. This highlights the role of the concept region mask in safeguarding each concept’s unique attributes. In the last column, concept features, such as the anime boy’s haircut being influenced by another character, and an unintended dog appearing, suggest leakage into unintended areas, due to down-sampling in U-Net. The inclusion of $\mathcal{L}_{region}$ addresses this issue by ensuring concept generation within specified regions only. These strategies validate the essential roles played by the Concept Enhancement and Concept Isolation Constraints in maintaining concept integrity and precision within the generated images, significantly bolstering the model’s capability to produce conceptually coherent and visually accurate outputs.
0.B.2 More Comparison
@@
Figure 9: Comparison with Mix-of-Show, where image-based conditions are applied. The yellow box emphasizes the issue of concept confusion, while red boxes underscore instances of concept vanishing.
@@
To facilitate a fair comparison with Mix-of-Show[7], we follow their default setting, applying identical image-based conditions for customizing multi-concept images using the same random seed. The results are illustrated in Fig. 9, where we display four distinct concept combinations in both anime and real-world styles. Our observations indicate that Mix-of-Show encounters challenges in maintaining identity features (indicated by yellow boxes) and achieving the completeness of concepts (highlighted by red boxes). However, our method effectively addresses these issues, delivering high-fidelity and coherent images to users, enhancing the overall quality and satisfaction with the generated content.
0.B.3 User Study
@@
Participants rated each aspect on a scale from 1 to 5, where higher scores denote superior quality. To thoroughly gauge the performance across various multi-object customization scenarios, we included setups involving 2, 3, 4, and 5 customization concepts. The sequence of all image-question pairs was randomized before being presented to 25 different users for evaluation. Each user was tasked with rating a total of 60 questions. The study results are shown in Tab. 3. Across all scenarios, LoRA-Composer emerged as the preferred choice, receiving the highest score of votes. Notably, our method demonstrated significant strengths, especially in scenarios that required eliminating image-based conditions. These outcomes demonstrate the effectiveness of LoRA-Composer in the generation of multi-concept customized images.
@@
0.B.4 More Visual Results
@@
Figure 10: More results of our method in three configurations.
Appendix 0.C Potential Negative Society Impact
@@
Appendix 0.D Limitation
@@
The second limitation concerns concepts beyond layout boundaries (Fig. 11(b)): Sometimes, foreground pixels might exceed their intended layout boundaries, a result of the model’s (Stable Diffusion’s [26]) design to yield results based on common sense. Opting for a more reasonable layout strategy could remedy this.
The final limitation concerns inference efficiency, where a notable delay occurs due to the requirement to load various LoRAs checkpoints and perform backward computations for updating latent representations.

### v2

@@ C.1 Default setting in experiments
C.2 More Ablation Study
@@
Figure 8: More ablation study on concept enhancement constraints and concept isolation constraints. The upper portion displays outcomes utilizing our full methodology, while the lower portion illustrates results with specific modules omitted, highlighting the significance of each component within our approach. The red boxes and the yellow boxes are used to accentuate the distinctions between the anime style and the real-world style, respectively.
@@
In the main ablation study (section 4.5), we explored the synergistic effects of combining modules with similar functionalities. Here, we delve deeper with an extensive ablation study on our concept enhancement constraints, which includes Gaussian sample strategy in eq. 3 and $\mathcal{L}_{f}$ in eq. 6) and concept isolation constraints (incorporating the concept region mask and $\mathcal{L}_{r}$ in eq. 7). These examinations aim to illuminate their contributions to model performance, as illustrated in Fig. 8. Specifically, in the first column, the absence of Gaussian sampling leads to the concepts not being accurately centered within their designated boxes. This lack of precision can even cause anime concepts to appear outside their intended boundaries, resulting in a loss of their unique identity traits. In the second column, without $\mathcal{L}_{f}$, both anime and realistic figures fail to occupy their designated boxes completely, pointing to a deficiency in fully utilizing the allocated space. In the third column, we observe concept confusion, characterized by the merging of anime haircuts and the loss of distinctive facial traits in realistic figures, which indicates a loss of distinctiveness. This highlights the role of the concept region mask in safeguarding each concept’s unique attributes. In the last column, concept features, such as the anime boy’s haircut being influenced by another character, and an unintended dog appearing. These issues indicate that there is leakage into unintended areas, due to down-sampling in U-Net. The inclusion of $\mathcal{L}_{r}$ effectively addresses this problem by minimizing the influence of background elements on foreground concepts. These strategies validate the essential roles played by the concept enhancement and concept isolation constraints in maintaining concept integrity and precision within the generated images, significantly bolstering the model’s capability to produce conceptually coherent and visually accurate outputs.
C.3 More Comparison
Figure 9: More qualitative comparison with baselines. For each case, we use the same seeds.
@@ C.3 More Comparison
Figure 10: Comparison with Mix-of-Show, where image-based conditions are applied. The yellow box emphasizes the issue of concept confusion, while the red boxes underscore instances of concept vanishing.
@@
To ensure a fair comparison with Mix-of-Show [2], we adopted their default settings, applying the same image-based conditions and using identical random seeds for generating multi-concept images. The comparative results are depicted in Fig. 10, showcasing four unique concept combinations styled in both anime and real-world visuals. Our analysis reveals that while Mix-of-Show struggles with maintaining distinct identity features (as indicated by yellow boxes) and the completeness of the integrated concepts (highlighted by red boxes), our approach successfully overcomes these limitations. Our method produces high-fidelity, coherent images that significantly enhance user satisfaction and improve the perceived quality of the generated content.
Figure 11: The user study interface that participants used to evaluate generated images on text-to-image alignment and image-to-image alignment.
C.4 User Study
@@
As shown in Fig. 11, participants rated each aspect on a scale from 1 to 5, where higher scores denote superior quality. To thoroughly gauge the performance across various multi-object customization scenarios, we included setups involving 2, 3, 4, and 5 customization concepts. The sequence of all image-question pairs was randomized before being presented to 25 different users for evaluation. Each user was tasked with rating a total of 60 questions. The study results are shown in Table 3. Across all scenarios, LoRA-Composer emerged as the preferred choice, receiving the highest score of votes. Notably, our method demonstrated significant strengths, especially in scenarios that required eliminating image-based conditions. These outcomes demonstrate the effectiveness of LoRA-Composer in the generation of multi-concept customized images.
@@
Figure 12: More results of our method in four configurations.
C.5 More Visual Results
@@
Appendix D Potential Negative Society Impact
@@
Appendix E Limitation and Future Work
@@
The second limitation pertains to instances where concepts extend beyond their designated layout boundaries, as shown in Fig. 13(b). Occasionally, foreground elements may spill over their intended borders, a consequence of Stable Diffusion’s [1] design, which relies on generalized assumptions to generate outcomes. Adopting a more structured layout strategy could potentially mitigate this issue.
The final limitation pertains to inference efficiency. A slight delay occurs due to the need to load various LoRA checkpoints and perform backward computations to update latent representations. This process takes approximately 20-40 seconds per image on a single NVIDIA A100 GPU.